# Compression for Adaptive Edge LLMs

### v2

@@ Abstract
00footnotetext: #  Corresponding authors.
@@ 1 Introduction
Numerous emerging applications, such as question answering on VR/AR glasses, are powered by large language models (LLMs). Yet, models with parameters on the order of billions (e.g., 10B) restrict the platforms and applications that can utilize them. Extensive research investigates quantization (xiao2023smoothquant; lin2024awq; lin2024qserve) and compression (qinsi2025dobisvd; wang2025svdllm; lin2025modegpt; wang2025bitstack) for LLMs to lower memory and computing needs for deployment. However, the limited and shared resources (e.g., the unified memory architecture) on edge devices still pose huge challenges for model deployment. Since resources (e.g., memory) are dynamically managed by the operating system, the availability of the resources highly depends on the system workload. As a result, the pre-compressed or pre-quantized language models with fixed model sizes may not run on a device under high workload scenarios.
@@
Re-compressing or re-quantizing the model to fit it into available memory is not practical due to the high computational costs, i.e., several hours on cloud GPUs (lin2025modegpt; frantar2023gptq). A solution to address this issue is storing several model replicas at different compression rates. Nonetheless, producing pre-compressed replicas of different sizes is both time- and storage-consuming. Alternatively, employing elastic training (cai2024flextron; cai2025llamaflex) to a pre-trained model enables the derivation of various sizes from the model. Yet, this approach requires availability of GPU resources and training on curated datasets to support flexible deployment for one specific type of model, e.g., Llama-3.1-8B, limiting the applicability.
@@
Our proposed work addresses this issue under the post-training setting when access to server-class GPUs and curated datasets is limited. As illustrated in Figure 1, our framework supports quantization and structured pruning, performing efficiently on one server GPU. Our objective is to support and design compression algorithms for various model architectures, including Transformers, State Space Models (SSMs), and hybrid models. Our pipeline is shown in Figure 2. We group the weights within the block, gather channel corrections from a calibration dataset, and apply weight-sorting algorithms. Our multi-layer perceptron (MLP) weights are decomposed without any gradient information or expensive full matrix pseudo-inverse, yielding a speedup of $20\times$ compared to prior art (lin2025modegpt). For $W_{v}$ and $W_{o}$ in self-attention layers, we develop a quantization-aware singular value decomposition (SVD) of weights to minimize quantization errors. For SSMs and hybrids models, we find that SSM blocks are particularly sensitive to state matrices, and propose a state-aware weight-sorting strategy to mitigate this. We then apply a masked fine-tuning to the sorted model. In each fine-tuning step, a global pruning rate $P_{t}$ is chosen randomly, masking the least ranked channels in the layers. The refined model is then quantized in low bit-width and deployed on the edge platform. The entire process is performed once in the cloud. For the deployed model, we prune the models according to a specified global pruning rate, e.g., $P_{35}=35\%$, on the edge device. Our contributions are summarized as follows:
@@ 2 Related work
Prior work has aimed to reduce the size of Transformer-based LLMs for efficient deployment by utilizing low bit-width data types (xiao2023smoothquant; lin2024qserve; lin2024awq; zhao2024atom; ashkboos2024quarot; liu2025spinquant), minimizing storage needs and optimizing hardware for low-bit computations. Unstructured (frantar2023sparsegpt; sun2024simple) and semi-structured pruning (e.g., N:M sparsity) (li2023sparse) for reducing model size by removing specific parameters while minimizing accuracy loss. Nonetheless, deploying such methods requires specialized hardware (taka2025systolic; xia2023flash). Structured pruning (wang2025svd; lin2025modegpt; ma2023llm; ashkboos2024slicegpt) removes whole elements (e.g., channels and heads), enabling faster inference on standard hardware but potentially decreasing performance. Some studies focus on one-shot compression (genzel2025compressing; wang2025bitstack), flexible bit-width quantization (park2024any), and quantization with semi-structured sparsity (mozaffari2025slim). wang2025bitstack propose a solution for any size compression in FP16 models by iteratively adding SVD residual terms, but this method incurs substantial overhead, leading to higher latency compared to the original FP16 model. Our framework is systematically designed for quantization and on-device structured pruning with practical speedups.
@@
State Space Models are memory-efficient alternatives to Transformers. Recent studies (xu2025mambaquant; chiang2025quamba; chiang2025quamba2) introduce low-bit quantization techniques for SSMs. Structured (taghibakhshi2025efficient; munoz2025mamba) and unstructured pruning (tuo2025sparsessm; shihab2025efficient) strategies have been developed for SSMs. For example, taghibakhshi2025efficient eliminate the SSM heads and restore performance through knowledge distillation training. munoz2025mamba explore block-wise (e.g., Mamba and Transformer blocks) and module-wise (e.g., SSM and self-attention modules) structured pruning methods. Some work explores the token pruning (zhan2024exploring) or dimension reduction (chi2024vmean) for vision SSMs. Our focus is on analyzing a broader structured pruning and quantization framework for Transformers, SSMs, and hybrids, distinguishing it from previously mentioned approaches.
@@
Elastic training aims to enable a pre-trained LLM to dynamically adapt to varying deployment constraints such as memory, compute, and latency budgets. Flextron (cai2024flextron) and LLaMaFlex (cai2025llamaflex) introduce many-in-one architectures through pruning and weight sharing, allowing adaptive inference under dynamic resource constraints. Jet-Nemotron (gu2025jet) further leverages post-training neural architecture search to generate compact LLM variants. These methods require GPU resources and training on curated datasets for flexible deployment tailored to a particular model type and size, e.g., Llama-3.1-8B, thereby restricting their general applicability. In contrast, our work focuses on post-training on a single server GPU for various architectures and supports on-device adaptation.
@@ 3.2 Structured Weight Sorting
Our objective is to enable adaptive on-device pruning by sorting the weights according to their importance scores, allowing the device to prune the least significant columns. Inspired by recent studies (lin2025modegpt; koike2025latentllm), we group the weights and conduct joint decomposition, as shown in Figure 3. We co-design the pruning algorithm alongside quantization and fused kernels for Transformers, SSMs, and hybrids.
@@
To derive $\mathbf{S}_{m}$ for sorting the weight matrices in the MLP layer, we collect the intermediate activation $\mathbf{X}_{\mathrm{int}}=\sigma\left(\mathbf{X}\mathbf{W}_{g}\right)\odot\mathbf{X}\mathbf{W}_{u}$ from the calibration set and calculate the channel correlation $\mathcal{C}=\mathbf{X}_{\mathrm{int}}^{\top}\mathbf{X}_{\mathrm{int}}\in\mathbb{R}^{D_{\mathrm{int}}\times D_{\mathrm{int}}}$. We average the correlation matrix over the calibration set, and compute the ridge leverage scores (mccurdy2018ridge) defined by $\mathrm{diag}\left(\mathcal{C}(\mathcal{C}+\lambda I)^{-1}\right)$. We set ridge lambda $\lambda=1$ in our experiments. We use the scores and create a column sorting matrix $\mathbf{S}_{m}\in\mathbb{R}^{D_{\mathrm{int}}\times D_{\mathrm{int}}}$ that reorders the output columns for $\mathbf{W}_{u}$ and $\mathbf{W}_{g}$ as $\mathbf{W}_{u}\mathbf{S}_{m}$ and $\mathbf{W}_{g}\mathbf{S}_{m}$, and the input rows of the $\mathbf{W}_{d}$ as $\mathbf{S}_{m}^{\top}\mathbf{W}_{d}$, as shown in Figure 3 (a). We show the process in Algorithm 1.
@@
Our approach does not rely on time-consuming pseudo-inverse to sort the MLP weight matrices. Although the pseudo-inverse (i.e., Moore-Penrose inverse (penrose1955generalized)) provides a theoretical bound in pruning errors (lin2025modegpt), it exhibits three major drawbacks: (1) Pseudo-inverse has a complexity of $O(n^{3})$ for a $n$-size squared matrix. This is particularly time-consuming when computing the pseudo-inverse of correlation matrices in MLP layers because $D_{\mathrm{int}}$ is a large number in most LLM designs, e.g.,, Llama-3-8B $D_{\mathrm{int}}=14336$. (2) Pseudo-inverse computation requires a high-precision FP64 to maintain numerical stability (lin2025modegpt),which demands substantial memory usage for full-precision weights. (3) Matrix inverse breaks the equivalence of pruned weights, resulting in $(\mathbf{W^{\prime}})^{\dagger}\neq(\mathbf{W}^{\dagger})^{\prime}$, requiring recomputation for different pruning rates. Here, $\mathbf{W}\in\mathbb{R}^{D\times D}$. $\mathbf{W}^{\prime}$ is a submatrix of $\mathbf{W}$, where $\mathbf{W}^{\prime}\in\mathbb{R}^{D\times D^{\prime}}$ with $D>D^{\prime}$. $\mathbf{W}^{\dagger}$ represents the inverse matrix. We show the latency of pseudo-inverse computation for FP64 matrices on A6000 in Table 1.
@@
where $\rho(\cdot)$ denotes RoPE (su2021roformer). The weights in MHSA are divided into two groups: $\{\mathbf{W}^{i}_{q},\mathbf{W}^{i}_{k}\}$ and $\{\mathbf{W}^{i}_{v},\mathbf{W}^{i}_{o}\}$.
@@
For the $\mathbf{W}^{i}_{v}$ and $\mathbf{W}^{i}_{o}$, we perform an activation-scaled SVD decomposition (wang2025svd; yuan2023asvd). With the input correlation matrix $\mathcal{C}=\mathbf{X}^{\top}\mathbf{X}$, we follow lin2025modegpt to perform joint decomposition by two consecutive SVD operations: $\mathcal{C}^{\frac{1}{2}}\mathbf{W}^{i}_{v}\mathbf{W}^{i}_{o}=\mathrm{SVD}(\mathcal{C}^{\frac{1}{2}}\mathbf{W}^{i}_{v})\mathbf{W}^{i}_{o}=\mathbf{U}_{v}\mathbf{\Sigma}_{v}\mathbf{V^{\top}}_{v}\mathbf{W}^{i}_{o}=\mathbf{U}_{v}\mathrm{SVD}(\mathbf{\Sigma}_{v}\mathbf{V^{\top}}_{v}\mathbf{W}^{i}_{o})=\mathbf{U}_{v}\mathbf{U}\mathbf{\Sigma}\mathbf{V^{\top}}$. The SVD decomposition ranks the eigenvectors by eigenvalues. To reduce quantization errors, we fuse the diagonal matrix $\mathbf{\Sigma}$ into $\mathbf{W}^{i}_{v}$, unlike prior art (wang2025svdllm; lin2025modegpt). We present the final sorted weights with quantization-aware SVD decomposition as $\mathbf{W}^{i}_{v}=\mathcal{C}^{\frac{-1}{2}}\mathbf{U}_{v}\mathbf{U}\mathbf{\Sigma}$ and $\mathbf{W}^{i}_{o}=\mathbf{V^{\top}}$, as shown in Figure 3 (b). Low-bit quantization (i.e., INT4) is sensitive to numerical distribution within the quantization group. We deconstruct the weight matrix $\mathbf{W}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}$ by merging the long-tailed eigenvalues $\mathbf{\Sigma}$ with $\mathbf{U}$, such that $\mathbf{W}=(\mathbf{U}\mathbf{\Sigma})\mathbf{V}$, where each column of $\mathbf{U}$ is scaled by its eigenvalue $\sigma_{i}$. Thus, $\sigma_{i}$ acts as the quantization scaling factor for the group without distorting the distributions, as depicted in Figure 4 (b).
We show the details of our algorithms in Algorithm 2 and 3. This joint weight decomposition also supports Grouped-Query Attention (GQA) (ainslie2023gqa), as shown in Figure 3 (b) and Algorithm 6 and 7 in Appendix A.
@@
where $\{\mathbf{W}^{i}_{x},\mathbf{W}^{i}_{z}\}\in\mathbb{R}^{D_{\mathrm{h}}\times D_{\mathrm{hd}}}$, and $\{\mathbf{W}^{g}_{B},\mathbf{W}^{g}_{C}\}\in\mathbb{R}^{D_{\mathrm{h}}\times D_{\mathrm{s}}}$, and the output weight $\mathbf{W}^{i}_{x}\in\mathbb{R}^{D_{\mathrm{hd}}\times D_{\mathrm{h}}}$. $\mathrm{Norm}(\cdot)$ and $\phi(\cdot)$ denote normalization and 1D causal convolution layer fused with an activation, respectively. The $\mathrm{SSM}(\cdot)$ function performs the linear recurrence computations $h_{t}=\Delta_{t}A_{t}h_{t-1}+\Delta_{t}B_{t}x_{t},\quad y_{t}=C_{t}h_{t}$ for the $i^{\mathrm{th}}$ head and $g^{\mathrm{th}}$ group at each time step $t$ with a parameterized step size $\Delta$ (dao2024transformers). $\Delta{\mathbf{A}}$ and $\Delta\mathbf{B}^{g}=\Delta\phi(\mathbf{X}\mathbf{W}^{g}_{B})$ are the matrix forms of $\Delta_{t}A_{t}$ and $\Delta_{t}B_{t}$. $\mathcal{H}$ is the matrix form of the SSM state $h_{t}$. To perform the joint weight decomposition, we first break the computation of a Mamba block into two sub-formulations: (1) SSM input mask $\mathcal{M}$: $f_{\mathcal{M}}(\mathbf{X},g)=\mathbf{C}^{g}(\Delta\mathbf{B}^{g})^{\top}=\phi(\mathbf{X}\mathbf{W}^{g}_{C})\;(\Delta\phi(\mathbf{X}\mathbf{W}^{g}_{B}))^{\top}$ and (2) SSM state $\mathcal{H}$: $f_{\mathcal{H}}(\mathbf{X},g,i,h_{0})=\Delta{\mathbf{A}}\mathcal{H}(h_{0})\;+\Delta\mathbf{B}^{g}\mathbf{X}^{i}_{\phi}$ with an initial state $h_{0}$. We denote $\mathbf{X}^{i}_{\phi}=\phi(\mathbf{X}\mathbf{W}^{i}_{x})$ as the output of the causal convolution activation.

### v3

@@ Abstract
00footnotetext:  Corresponding authors.
@@ 1 Introduction
Numerous emerging applications, such as question answering on VR/AR glasses, are powered by large language models (LLMs). Yet, models with parameters on the order of billions (e.g., 10B) restrict the platforms and applications that can utilize them. Extensive research investigates quantization (Xiao et al., 2023; Lin et al., 2024a, b) and compression (Qinsi et al., 2025; Wang et al., 2025b; Lin et al., 2025; Wang et al., 2025c) for LLMs to lower memory and computing needs for deployment. However, the limited and shared resources (e.g., the unified memory architecture) on edge devices still pose huge challenges for model deployment. Since resources (e.g., memory) are dynamically managed by the operating system, the availability of the resources highly depends on the system workload. As a result, the pre-compressed or pre-quantized language models with fixed model sizes may not run on a device under high workload scenarios.
@@
Re-compressing or re-quantizing the model to fit it into available memory is not practical due to the high computational costs, i.e., several hours on cloud GPUs (Lin et al., 2025; Frantar et al., 2023). A solution to address this issue is storing several model replicas at different compression rates. Nonetheless, producing pre-compressed replicas of different sizes is both time- and storage-consuming. Alternatively, employing elastic training (Cai et al., 2024, 2025) to a pre-trained model enables the derivation of various sizes from the model. Yet, this approach requires availability of GPU resources and training on curated datasets to support flexible deployment for one specific type of model, e.g., Llama-3.1-8B, limiting the applicability.
@@
Our proposed work addresses this issue under the post-training setting when access to server-class GPUs and curated datasets is limited. As illustrated in Figure 1, our framework supports quantization and structured pruning, performing efficiently on one server GPU. Our objective is to support and design compression algorithms for various model architectures, including Transformers, State Space Models (SSMs), and hybrid models. Our pipeline is shown in Figure 2. We group the weights within the block, gather channel corrections from a calibration dataset, and apply weight-sorting algorithms. Our multi-layer perceptron (MLP) weights are decomposed without any gradient information or expensive full matrix pseudo-inverse, yielding a speedup of $20\times$ compared to prior art (Lin et al., 2025). For $W_{v}$ and $W_{o}$ in self-attention layers, we develop a quantization-aware singular value decomposition (SVD) of weights to minimize quantization errors. For SSMs and hybrids models, we find that SSM blocks are particularly sensitive to state matrices, and propose a state-aware weight-sorting strategy to mitigate this. We then apply a masked fine-tuning to the sorted model. In each fine-tuning step, a global pruning rate $P_{t}$ is chosen randomly, masking the least ranked channels in the layers. The refined model is then quantized in low bit-width and deployed on the edge platform. The entire process is performed once in the cloud. For the deployed model, we prune the models according to a specified global pruning rate, e.g., $P_{35}=35\%$, on the edge device. Our contributions are summarized as follows:
@@ 2 Related work
Prior work has aimed to reduce the size of Transformer-based LLMs for efficient deployment by utilizing low bit-width data types (Xiao et al., 2023; Lin et al., 2024b, a; Zhao et al., 2024; Ashkboos et al., 2024b; Liu et al., 2025), minimizing storage needs and optimizing hardware for low-bit computations. Unstructured (Frantar and Alistarh, 2023; Sun et al., 2024) and semi-structured pruning (e.g., N:M sparsity) (Li et al., 2023) for reducing model size by removing specific parameters while minimizing accuracy loss. Nonetheless, deploying such methods requires specialized hardware (Taka et al., 2025; Xia et al., 2023). Structured pruning (Wang et al., 2025a; Lin et al., 2025; Ma et al., 2023; Ashkboos et al., 2024a) removes whole elements (e.g., channels and heads), enabling faster inference on standard hardware but potentially decreasing performance. Some studies focus on one-shot compression (Genzel et al., 2025; Wang et al., 2025c), flexible bit-width quantization (Park et al., 2024), and quantization with semi-structured sparsity (Mozaffari et al., 2025). Wang et al. (2025c) propose a solution for any size compression in FP16 models by iteratively adding SVD residual terms, but this method incurs substantial overhead, leading to higher latency compared to the original FP16 model. Our framework is systematically designed for quantization and on-device structured pruning with practical speedups.
@@
State Space Models are memory-efficient alternatives to Transformers. Recent studies (Xu et al., 2025; Chiang et al., 2025b, a) introduce low-bit quantization techniques for SSMs. Structured (Taghibakhshi et al., 2025; Muñoz et al., 2025) and unstructured pruning (Tuo and Wang, 2025; Shihab et al., 2025) strategies have been developed for SSMs. For example, Taghibakhshi et al. (2025) eliminate the SSM heads and restore performance through knowledge distillation training. Muñoz et al. (2025) explore block-wise (e.g., Mamba and Transformer blocks) and module-wise (e.g., SSM and self-attention modules) structured pruning methods. Some work explores the token pruning (Zhan et al., 2024) or dimension reduction (Chi et al., 2024) for vision SSMs. Our focus is on analyzing a broader structured pruning and quantization framework for Transformers, SSMs, and hybrids, distinguishing it from previously mentioned approaches.
@@
Elastic training aims to enable a pre-trained LLM to dynamically adapt to varying deployment constraints such as memory, compute, and latency budgets. Flextron (Cai et al., 2024) and LLaMaFlex (Cai et al., 2025) introduce many-in-one architectures through pruning and weight sharing, allowing adaptive inference under dynamic resource constraints. Jet-Nemotron (Gu et al., 2025) further leverages post-training neural architecture search to generate compact LLM variants. These methods require GPU resources and training on curated datasets for flexible deployment tailored to a particular model type and size, e.g., Llama-3.1-8B, thereby restricting their general applicability. In contrast, our work focuses on post-training on a single server GPU for various architectures and supports on-device adaptation.
@@ 3.2 Structured Weight Sorting
Our objective is to enable adaptive on-device pruning by sorting the weights according to their importance scores, allowing the device to prune the least significant columns. Inspired by recent studies (Lin et al., 2025; Koike-Akino et al., 2025), we group the weights and conduct joint decomposition, as shown in Figure 3. We co-design the pruning algorithm alongside quantization and fused kernels for Transformers, SSMs, and hybrids.
@@
To derive $\mathbf{S}_{m}$ for sorting the weight matrices in the MLP layer, we collect the intermediate activation $\mathbf{X}_{\mathrm{int}}=\sigma\left(\mathbf{X}\mathbf{W}_{g}\right)\odot\mathbf{X}\mathbf{W}_{u}$ from the calibration set and calculate the channel correlation $\mathcal{C}=\mathbf{X}_{\mathrm{int}}^{\top}\mathbf{X}_{\mathrm{int}}\in\mathbb{R}^{D_{\mathrm{int}}\times D_{\mathrm{int}}}$. We average the correlation matrix over the calibration set, and compute the ridge leverage scores (McCurdy, 2018) defined by $\mathrm{diag}\left(\mathcal{C}(\mathcal{C}+\lambda I)^{-1}\right)$. We set ridge lambda $\lambda=1$ in our experiments. We use the scores and create a column sorting matrix $\mathbf{S}_{m}\in\mathbb{R}^{D_{\mathrm{int}}\times D_{\mathrm{int}}}$ that reorders the output columns for $\mathbf{W}_{u}$ and $\mathbf{W}_{g}$ as $\mathbf{W}_{u}\mathbf{S}_{m}$ and $\mathbf{W}_{g}\mathbf{S}_{m}$, and the input rows of the $\mathbf{W}_{d}$ as $\mathbf{S}_{m}^{\top}\mathbf{W}_{d}$, as shown in Figure 3 (a). We show the process in Algorithm 1.
@@
Our approach does not rely on time-consuming pseudo-inverse to sort the MLP weight matrices. Although the pseudo-inverse (i.e., Moore-Penrose inverse (Penrose, 1955)) provides a theoretical bound in pruning errors (Lin et al., 2025), it exhibits three major drawbacks: (1) Pseudo-inverse has a complexity of $O(n^{3})$ for a $n$-size squared matrix. This is particularly time-consuming when computing the pseudo-inverse of correlation matrices in MLP layers because $D_{\mathrm{int}}$ is a large number in most LLM designs, e.g.,, Llama-3-8B $D_{\mathrm{int}}=14336$. (2) Pseudo-inverse computation requires a high-precision FP64 to maintain numerical stability (Lin et al., 2025),which demands substantial memory usage for full-precision weights. (3) Matrix inverse breaks the equivalence of pruned weights, resulting in $(\mathbf{W^{\prime}})^{\dagger}\neq(\mathbf{W}^{\dagger})^{\prime}$, requiring recomputation for different pruning rates. Here, $\mathbf{W}\in\mathbb{R}^{D\times D}$. $\mathbf{W}^{\prime}$ is a submatrix of $\mathbf{W}$, where $\mathbf{W}^{\prime}\in\mathbb{R}^{D\times D^{\prime}}$ with $D>D^{\prime}$. $\mathbf{W}^{\dagger}$ represents the inverse matrix. We show the latency of pseudo-inverse computation for FP64 matrices on A6000 in Table 1.
@@
where $\rho(\cdot)$ denotes RoPE (Su et al., 2021). The weights in MHSA are divided into two groups: $\{\mathbf{W}^{i}_{q},\mathbf{W}^{i}_{k}\}$ and $\{\mathbf{W}^{i}_{v},\mathbf{W}^{i}_{o}\}$.
@@
For the $\mathbf{W}^{i}_{v}$ and $\mathbf{W}^{i}_{o}$, we perform an activation-scaled SVD decomposition (Wang et al., 2025a; Yuan et al., 2023). With the input correlation matrix $\mathcal{C}=\mathbf{X}^{\top}\mathbf{X}$, we follow Lin et al. (2025) to perform joint decomposition by two consecutive SVD operations: $\mathcal{C}^{\frac{1}{2}}\mathbf{W}^{i}_{v}\mathbf{W}^{i}_{o}=\mathrm{SVD}(\mathcal{C}^{\frac{1}{2}}\mathbf{W}^{i}_{v})\mathbf{W}^{i}_{o}=\mathbf{U}_{v}\mathbf{\Sigma}_{v}\mathbf{V^{\top}}_{v}\mathbf{W}^{i}_{o}=\mathbf{U}_{v}\mathrm{SVD}(\mathbf{\Sigma}_{v}\mathbf{V^{\top}}_{v}\mathbf{W}^{i}_{o})=\mathbf{U}_{v}\mathbf{U}\mathbf{\Sigma}\mathbf{V^{\top}}$. The SVD decomposition ranks the eigenvectors by eigenvalues. To reduce quantization errors, we fuse the diagonal matrix $\mathbf{\Sigma}$ into $\mathbf{W}^{i}_{v}$, unlike prior art (Wang et al., 2025b; Lin et al., 2025). We present the final sorted weights with quantization-aware SVD decomposition as $\mathbf{W}^{i}_{v}=\mathcal{C}^{\frac{-1}{2}}\mathbf{U}_{v}\mathbf{U}\mathbf{\Sigma}$ and $\mathbf{W}^{i}_{o}=\mathbf{V^{\top}}$, as shown in Figure 3 (b). Low-bit quantization (i.e., INT4) is sensitive to numerical distribution within the quantization group. We deconstruct the weight matrix $\mathbf{W}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}$ by merging the long-tailed eigenvalues $\mathbf{\Sigma}$ with $\mathbf{U}$, such that $\mathbf{W}=(\mathbf{U}\mathbf{\Sigma})\mathbf{V}$, where each column of $\mathbf{U}$ is scaled by its eigenvalue $\sigma_{i}$. Thus, $\sigma_{i}$ acts as the quantization scaling factor for the group without distorting the distributions, as depicted in Figure 4 (b).
We show the details of our algorithms in Algorithm 2 and 3. This joint weight decomposition also supports Grouped-Query Attention (GQA) (Ainslie et al., 2023), as shown in Figure 3 (b) and Algorithm 6 and 7 in Appendix A.
@@
where $\{\mathbf{W}^{i}_{x},\mathbf{W}^{i}_{z}\}\in\mathbb{R}^{D_{\mathrm{h}}\times D_{\mathrm{hd}}}$, and $\{\mathbf{W}^{g}_{B},\mathbf{W}^{g}_{C}\}\in\mathbb{R}^{D_{\mathrm{h}}\times D_{\mathrm{s}}}$, and the output weight $\mathbf{W}^{i}_{x}\in\mathbb{R}^{D_{\mathrm{hd}}\times D_{\mathrm{h}}}$. $\mathrm{Norm}(\cdot)$ and $\phi(\cdot)$ denote normalization and 1D causal convolution layer fused with an activation, respectively. The $\mathrm{SSM}(\cdot)$ function performs the linear recurrence computations $h_{t}=\Delta_{t}A_{t}h_{t-1}+\Delta_{t}B_{t}x_{t},\quad y_{t}=C_{t}h_{t}$ for the $i^{\mathrm{th}}$ head and $g^{\mathrm{th}}$ group at each time step $t$ with a parameterized step size $\Delta$ (Dao and Gu, 2024). $\Delta{\mathbf{A}}$ and $\Delta\mathbf{B}^{g}=\Delta\phi(\mathbf{X}\mathbf{W}^{g}_{B})$ are the matrix forms of $\Delta_{t}A_{t}$ and $\Delta_{t}B_{t}$. $\mathcal{H}$ is the matrix form of the SSM state $h_{t}$. To perform the joint weight decomposition, we first break the computation of a Mamba block into two sub-formulations: (1) SSM input mask $\mathcal{M}$: $f_{\mathcal{M}}(\mathbf{X},g)=\mathbf{C}^{g}(\Delta\mathbf{B}^{g})^{\top}=\phi(\mathbf{X}\mathbf{W}^{g}_{C})\;(\Delta\phi(\mathbf{X}\mathbf{W}^{g}_{B}))^{\top}$ and (2) SSM state $\mathcal{H}$: $f_{\mathcal{H}}(\mathbf{X},g,i,h_{0})=\Delta{\mathbf{A}}\mathcal{H}(h_{0})\;+\Delta\mathbf{B}^{g}\mathbf{X}^{i}_{\phi}$ with an initial state $h_{0}$. We denote $\mathbf{X}^{i}_{\phi}=\phi(\mathbf{X}\mathbf{W}^{i}_{x})$ as the output of the causal convolution activation.